OpenStudy (anonymous):

i already asked how to find the vertex but know i need to find the x and y intercepts to x^2+8x+7

OpenStudy (anonymous):

set the equation equal to zero and you will find the x intercept(s)

OpenStudy (anonymous):

unfoil the equation

OpenStudy (anonymous):

(x+1)(x+7)=0

OpenStudy (anonymous):

solve for x so x+1=0 and x+7=0

OpenStudy (anonymous):

-7 -1

OpenStudy (anonymous):

then x= -1 and x= -7

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

is that all you have to do to solve for x and y?

OpenStudy (anonymous):

for the x intercept yes, but the y intercept you set the x values to zero and solve

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

0^2 + 8(0) + 7 = y

OpenStudy (anonymous):

thus y=7 and this is your y-intercept